Circular No. 8422

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN THE LARGE MAGELLANIC CLOUD W. Liller, Viña del Mar, Chile, reports his discovery of a possible nova

(V ≃ 10.8) near the position of the presumed 1937 nova YY Dor (without providing a measured position for the new object) on two unfiltered Technical Pan photographs taken on Oct. 20.193 UT; nothing was present at this location on his 0.2-m Schmidt camera photos taken on Oct. 17.181. A visual magnitude estimate by A. Pearce (Nedlands, W. Australia) on Oct. 21.576 gave 12.0. L. A. G. Monard, Pretoria, South Africa, reports that a CCD image taken on Oct. 21.88 shows a star of mag V ∼ 12.0 located at α = 5^{h}56^{m}41^{s}.84, δ = −68^{o}54′37′′.1 (equinox 2000.0), adding that a USNO-A2.0 star with blue mag 18.8 and red mag 17.8 has position end figures 42^{s}.38, 34′′.3. This appears to be some distance from YY Dor [cf. IBVS 1248, which cited Hoffmeister 1949, Astron. Abhandl. 12(1), A24], for which an accurate position is not available (the high precision given by Subramaniam and Anupama 2001, A.Ap. 449, 451, seems unwarranted). Henize et al. (1954, Proc. Nat. Acad. Sci. 40, 368) gave α = 5^{h}57^{m}.0, δ = −68^{o}55′ (equinox 2000.0, precessed from the original 1900.0 coordinates; earlier raw X,Y plate coordinates provided by McKibben 1941, Harv. Bull.

915) from Harvard/Boyden plates for the 1937 nova, and Hoffmeister gave α = 5^{h}57^{m}30^{s}, δ = −68^{o}55′.1 (equinox 2000.0; precessed from the original 1875.0 coordinates) from Windhoek plates.

SUPERNOVAE 2004ex AND 2004fc Further to IAUC 8412, K. Shimasaki and W. Li report the LOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Oct. 21.39 (mag ∼ 15.5) and 22.40 UT (mag ∼ 15.4). SN 2004fc is located at α = 1^{h}51^{m}03^{s}.85, δ = −9^{o}42′06′′.9 (equinox 2000.0), which is ∼ 2′′ east and 2′′ north of the diffuse nucleus of NGC 701. A KAIT image taken on Oct. 14.39 (limiting mag ∼ 19.0) showed nothing at this position. H. Yamaoka, Kyushu University, reports the following prediscovery unfiltered CCD magnitudes of SN 2004fc by K. Itagaki (Teppo-cho, Yamagata, Japan, 0.60-m reflector): Oct. 16.624, 15.4; 17.577, 15.4; it was not detected on an image taken on Oct. 6.718 (limiting magnitude difficult to estimate due to proximity to host-galaxy nucleus).

An independent discovery of SN 2004ex (cf. IAUC 8418, 8420) was made by Shimasaki and Li, with the following approximate unfiltered CCD magnitudes available: Oct. 6.35, [20.0; 10.33, [19.0; 14.33, 18.6.

2004 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green